The area of a rectangular window is to 306 square centimeters. If the length exceeds the width by 1 centimeters, what are the dimentions?
How do I solve this problem? I want to learn step by step.

Answer by mananth(16946)

The area of a rectangular window is to 306 square centimeters. If the length exceeds the width by 1 centimeters, what are the dimentions?
How do I solve this problem? I want to learn step by step.
let width be x
length = (x+1)
Area = 306
..
Area of rectangle A = L*W
A= x(x+1)= 306
x^2+x=306
add -306 to both sides
x^2+x-306 = 0
find factors of 306 such that their product = 306 and their sum is 1
x^2+18x-17x-306=0
x(x+18)-17(x+18)=0
(x-17)(x+18)=0
the width can be 17 OR -18. width cannot be -18. so ignore.
width = 17 cm
length = x+1 = 17+1=18 cm
..
the dimensions are 17 by 18 cm
...
m.ananth@hotmail.ca